solve using elimination.
8x-3y=-11
2x-5y=27

Answer by ewatrrr(24785)


Hi
solving using elimination.
8x-3y=-11
2x-5y=27
8x- 3y =-11
-8x+20y=27*-4 Mulitplying each term by -4 to eliminate x by adding the equations
17y = -119
y = -7
solving for x
8x-3*-7 =-11
8x +21 = -11
8x = -32
x = -4